OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are 20 and 2500, respectively.

OpenStudy (anonymous):

an = 4 • 5n + 1 an = 4 • 5n - 1 an = 4 • 5n + 3 an = 4 • 5n - 2

OpenStudy (anonymous):

would my answer be (c) is this right?

OpenStudy (anonymous):

@abb0t @Baby_Blount @Candy13106 @CaitlinN.

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

correct answer is an = 4 • 5^n - 1